An investment of $6,599.20 earns 4.2% interest compounded monthly over 7 years. Approximately how much interest is earned on the investment?
Answer by addingup(3677)

Amount = Principal(1+r/n)^nt
A = 6,599.20(1+0.042/12)^(12*7) = 8,850.16
8,850.16-6,599.20 = 2,250.96 is the amount interest earned on the investment